# Harmony in Divergence: Towards Fast, Accurate, and Memory-efficient Zeroth-order LLM Fine-tuning

### v1

@@ 6 Conclusion
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Related Work
A.1 Fine-tuning of Pre-trained Models
Fine-tuning a pre-trained model offers a powerful way to reuse learned representations and reduce training costs compared to building models from scratch, often achieving superior performance (Gururangan et al., 2020; Ouyang et al., 2022). Initially successful in NLP with models like BERT, RoBERTa, and GPT (Devlin, 2018; Liu et al., 2019; Chen et al., 2022), fine-tuning has also shown promise in vision tasks such as CLIP and SWAG (Radford et al., 2021; Singh et al., 2022). Recent parameter-efficient fine-tuning (PEFT), including LoRA (Hu et al., 2021), and prefix tuning (Li & Liang, 2021), further minimize resource needs by updating only a small subset of parameters, preserving most of the pre-trained weights and ensuring valuable knowledge is retained.
@@ A.2 Zeroth-order Optimization and Acceleration
ZO optimization emerges as an attractive technique that optimizes the model without backpropagation (Chen et al., 2023, 2017; Ye et al., 2018; Verma et al., 2023; Dhurandhar et al., 2018, 2019). Unlike most frequently used FO optimization which directly obtains and leverages the gradient for optimization, the zeroth-order method utilizes objective function value oracle only, estimating the gradient by finite differences. ZO method has a wide range of applications in machine learning fields, including adversarial attack and defense (Chen et al., 2017; Ye et al., 2018; Verma et al., 2023), machine learning explainability (Dhurandhar et al., 2018, 2019), reinforcement learning (Vemula et al., 2019), and on-chip training (Gu et al., 2021). Recently, the ZO method has been proposed to be leveraged on LLM fine-tuning to address the significant memory usage. Malladi et al. (2023) proposed MeZO, first scaling ZO optimization to fine-tuning parameter-intensive LLMs, greatly reducing memory utilization. On top of MeZO, Zhao et al. (2024) proposed HiZOO, leveraging the estimated Hessian information for better learning capacity, but reducing the throughput of MeZO to some extent.
ZO optimization, although it significantly saves memory, converges more slowly than FO methods due to higher variance from random search. Liu et al. (2018) introduced ZO-SVRG by incorporating variance reduction techniques (Johnson & Zhang, 2013). Shu et al. (2023) proposed using a Gaussian process to model objective function queries, thereby reducing query complexity and allowing more frequent queries to lower gradient variance. Sener & Koltun (2020) performed random search on a learned low-dimensional manifold, reducing the number of needed objective queries. However, existing ZO accelerators face two main challenges when adapting to ZO fine-tuning for LLMs. First, these approaches were typically designed for smaller-scale tasks involving fewer parameters and less data, and cannot be directly extended to large-scale LLMs. Second, many prior methods focus on improving query efficiency, whereas recent work has shown that a single query can suffice for LLM fine-tuning (Malladi et al., 2023). How to effectively accelerate ZO optimization on large model fine-tuning remains a problem.
@@ B.1 Datasets and Evaluation
Table 7: The hyperparameter for experiments. For DiZO and DiZO LoRA, we only show the setting of extra hyperparameters, and have the same setting in other common hyperparameters with MeZO and MeZO LoRA respectively.
@@
For the RoBERTa-large model, we use the following classification datasets: SST-2 (Socher et al., 2013), SST-5 (Socher et al., 2013), SNLI (Bowman et al., 2015), TREC (Voorhees & Tice, 2000), MNLI (Yao et al., 2020), and RTE (Dagan et al., 2005; Bar-Haim et al., 2006; Bentivogli et al., 2009; Giampiccolo et al., 2007). Following previous studies, we cap the test set size at 1000 samples. Two training settings are considered: $k=16$ and $k=512$, where we randomly select 16 or 512 samples per class for both training and validation.
For the OPT and Llama series models, we use the SuperGLUE benchmark (Wang et al., 2019), which includes RTE (Dagan et al., 2005; Bar-Haim et al., 2006; Bentivogli et al., 2009; Giampiccolo et al., 2007), CB (De Marneffe et al., 2019), BoolQ (Clark et al., 2019), WIC (Pilehvar & Camacho-Collados, 2018), WSC (Levesque et al., 2012), and MultiRC (Khashabi et al., 2018). We also include SST-2 (Socher et al., 2013) and two question answering datasets, SQuAD (Rajpurkar, 2016) and DROP (Dua et al., 2019). For each of these datasets, we randomly sample 1000 instances for training, 500 for validation, and 1000 for testing.
B.2 Implementation of Baselines
Memory-efficient ZO (MeZO) MeZO (Malladi et al., 2023) serves as a fundamental baseline for fine-tuning large language models (LLMs) using zeroth-order (ZO) optimization. By resampling perturbations with a fixed random seed, MeZO eliminates the need to store perturbations that are the same size as the model, thereby saving memory. For our implementation of MeZO, we adapted the code released by the authors at https://github.com/princeton-nlp/MeZO with minimal modifications.
Hessian-informed ZO (HiZOO) HiZOO (Zhao et al., 2024) is a recently proposed method for ZO fine-tuning of LLMs that leverages estimated second-order information to accelerate optimization. During the implementation of HiZOO, we identified several bugs in the released code at https://anonymous.4open.science/r/HiZOO-27F8, such as overflows when computing the Hessian. Consequently, we implemented the baseline ourselves. Additionally, we used the parameter settings from the original code instead of those described in the paper, as they resulted in better performance according to our implementation.
B.3 Hyperparameter Setting
We use the hyperparameters in Table 7 for experiments on RoBERTa-large, OPT-series, and Llama-series models. Specifically, the choice of clip range did not significantly impact the performance. The selection of the projection update cycle and scalar for projection affects the performance somewhat. Generally, for datasets that need larger iterations for convergence, or for these harder datasets, DiZO prefers a larger update cycle, while for those less complicated datasets, DiZO benefits from a smaller update cycle.
Appendix C Closer look at DiZO
@@ C.1 Ablation for Projection Layers Selection
Instead of applying projections to all layers, which would require storing the entire pre-trained model, we focus only on projecting the weights of the Query and Value in the attention modules. As shown in Table 8, this strategy achieves the best trade-off between the overall performance and extra storage requirements, does not reduce the performance and only 16.7% of the parameters of the pre-trained model are needed to store. A Similar strategy has also been adopted in LoRA (Hu et al., 2021).
Table 8: Ablation study for selecting which layers to project. The highlighted line with a blue rectangle is the setting used in DiZO. Extra memory indicates the extra memory needed due to pre-trained model storing. Attn_Q: attention Query layer; Attn_V: attention Value layer; Attn_K: attention Key layer; Attn_O: attention output projection; Dense: dense fully connected layer.
@@ C.2 Ablation for Strategies in ZO Projection Learning
As discussed in Section 3.3, we introduce two strategies, re-initialization (Re-init) and projection clipping (Clipping), to enhance projection learning and improve the stability of fine-tuning. The ablation results for these strategies, along with the corresponding loss curves, are shown in Figure 5.
Overall (left in Figure 5), omitting either Re-init or Clipping significantly diminishes the benefits of DiZO, with MeZO outperforming DiZO in these cases. Specifically, without Re-init, accuracy drops sharply, falling below MeZO. Similarly, without Clipping, while DiZO slightly outperforms MeZO on simpler datasets like SST-2, it suffers from severe model collapse on more challenging datasets, leading to a significant decline in accuracy.
From the loss curve trajectory (right in Figure 5), without Re-init, DiZO loses its advantage in training acceleration, as the loss curve becomes noticeably slower to decrease. Without Clipping, the loss curve exhibits significant oscillations during certain training steps. This instability arises when projections are optimized to unsuitable values, such as extremely large or small magnitudes. These inappropriate projections cause substantial changes in model weights, leading to pronounced oscillations in the loss.
@@
Figure 5: Ablation study for the two strategies: re-initialization and projection clipping, which is conducted on RoBERTa-large ($k=16$). Left: overall results when ablating the strategies. Right: loss curve when ablating the strategies.
C.3 Does Other Alternative Strategies for Layer-wise Divergence Work?
As discussed in Section 2.2, our objective is to enhance layer-wise divergence in ZO optimization. Naturally, with consideration of this objective, one may raise two questions regarding the projection strategy we adopt: 1) Can we perform layer-wise projections on the learning rate? 2) When updating weight by projection at $t$-th iteration, why do we use the weights of pre-trained model $\bm{\theta}_{0}$ as the base of the update (shown in Eq.5) instead of the weights from the $(t-1)$-th iteration, $\bm{\theta}_{t-1}$?
Table 9: Comparison on conducting projection on learning rate (LR) or use weight at $(t-1)$-th iteration $\bm{\theta}_{t-1}$ instead of the weight of the pre-trained model $\bm{\theta_{0}}$ as the base of projection. Results are obtained by fine-tuning OPT-2.7B.
Dataset SST-2 RTE SQuAD
@@ C.3 Does Other Alternative Strategies for Layer-wise Divergence Work?
To answer the above two questions. We investigate two alternative projection strategies: 1) searching layer-wise ideal learning rate via ZO optimization and then applying the updates, and 2) conducting projection update based on $\bm{\theta}_{t-1}$ instead of $\bm{\theta}_{0}$. Results are illustrated in Table 9, neither approach achieves performance comparable to DiZO; both yield results closer to MeZO in terms of accuracy and required GPU hours.
@@
Table 10: Memory utilization and speed test on OPT-2.7B on SST-2 dataset (35 tokens per example on average). : partial gradient-free; : gradient-free; : gradient-based. DiZO: searching projection with Adam.
@@
Table 11: Memory utilization and speed test on OPT-2.7B on SQuAD dataset (300 tokens per example on average). : partial gradient-free; : gradient-free; : gradient-based. DiZO: searching projection with Adam.
@@
We attribute this phenomenon to the high noise inherent in each ZO update, which relies on random perturbations and thus produces a highly imprecise update direction. In contrast, DiZO projects the optimization direction between the pre-trained and fine-tuned models, and this direction is supposed to be correct. Otherwise, the entire optimization would fail and the loss would not decrease. Moreover, recent studies suggest that the fine-tuned model is often less robust than their pre-trained version due to catastrophic forgetting (Dong et al., 2021; Oh et al., 2023; Zhai et al., 2023; Wang et al., 2024). Maintaining a connection with the pre-trained model helps robustify the fine-tuning process and mitigate some of the noise introduced by ZO’s random perturbations.
Appendix D More Experiment Results
D.1 Memory and Speed Analysis
We present the memory and speed results for OPT-2.7B on the SST-2 and SQuAD datasets in Table 10 and Table 11, respectively. DiZO significantly reduces the number of required iterations while maintaining throughput comparable to MeZO, leading to substantially fewer training GPU hours. In contrast, HiZOO achieves only modest iteration savings and further reduces the throughput of MeZO by approximately 1.5× due to its reliance on second-order information estimation. As a result, HiZOO offers only a slight improvement over MeZO in terms of training GPU hours. In some cases, such as HiZOO combined with LoRA on SQuAD, it even consumes more training GPU hours than MeZO with LoRA.
D.2 Llama Experiments
To demonstrate the broader applicability of DiZO, we conducted experiments on the Llama-series models. The results for Llama3-3B and Llama3-8B are presented in Table 12 and Table 13, respectively. DiZO consistently outperforms MeZO across both the 3B and 8B Llama models.
However, we observed that ZO LoRA performs poorly with Llama models (including DiZO, MeZO and HiZOO). The loss value remains stagnant, and the resulting accuracy is comparable to or even worse than zero-shot results. We leave it to future work to investigate why ZO LoRA fails with Llama models. We suspect that this limitation may be related to the Group Query Attention (GQA) (Ainslie et al., 2023) mechanism employed in Llama3.
Table 12: Experiments results on Llama3-3B for seven classification datasets and two text generation datasets (with 1000 training samples). Better results between MeZO and DiZO are highlighted in bold.

### v3

@@ Appendix A Related Work
A.1 Memory-efficient Fine-tuning
Fine-tuning a pre-trained model offers a powerful way to reuse learned representations and reduce training costs compared to building models from scratch, often achieving superior performance [36, 37]. Initially successful in NLP with models like BERT, RoBERTa, and GPT [38, 2, 39], fine-tuning has also shown promise in vision tasks such as CLIP and SWAG [40, 41].
Despite the success of fine-tuning, its high cost makes it not feasible. Therefore memory efficient fine-tuning method come up. Recent parameter-efficient fine-tuning (PEFT), including LoRA [17], and prefix tuning [42], minimize resource needs by updating only a small subset of parameters, preserving most of the pre-trained weights and ensuring valuable knowledge is retained. Low-rank decomposition-based methods, led by LoRA, have achieved remarkable success. The main idea is to minimize resource needs by updating only a small subset of parameters, preserving most of the pre-trained weights and ensuring valuable knowledge is retained. DoRA [43] decomposes the pre-trained weight into two components, magnitude and direction, to enhance both the learning capacity and training stability of LoRA. GaLore [44] proposed gradient low-rank projection, allows full-parameter learning while retaining the memory advantages of low-rank training. Beside low-rank method, quantization stood out as a promising method to reduce resources utilization. GPT3.int8() [45] identified the outlier in activation, and include a new mixed-precision decomposition scheme, which isolates the outlier feature dimensions into a 16-bit matrix multiplication while still more than 99.9% of values are multiplied in 8-bit. Despite training with mix-precision, SmoothQuant [46] smooths the activation outliers by offline by scaling, migrating the quantization difficulty from activations to weights. Moreover, Outlier Suppression+ [47] and OmniQuant [48] apply channel-wise shifting for asymmetry and channel-wise scaling for concentration for int4-level quantization.
@@ A.2 Zeroth-order Optimization and Acceleration
ZO optimization emerges as an attractive technique that optimizes the model without backpropagation [22, 49, 24, 18, 50, 19]. Unlike most frequently used FO optimization, which directly obtains and leverages the gradient for optimization, the zeroth-order method utilizes the objective function value oracle only, estimating the gradient by finite differences. ZO method has a wide range of applications in machine learning fields, including adversarial attack and defense [49, 24, 18], machine learning explainability [50, 19], reinforcement learning [51], and on-chip training [21]. Recently, the ZO method has been proposed to be leveraged on LLM fine-tuning to address the significant memory usage. [14] proposed MeZO, first scaling ZO optimization to fine-tuning parameter-intensive LLMs, greatly reducing memory utilization. On top of MeZO, [15] proposed HiZOO, leveraging the estimated Hessian information for better learning capacity, but reducing the throughput of MeZO to some extent.
ZO optimization, although it significantly saves memory, converges more slowly than FO methods due to higher variance from random search. [23] introduced ZO-SVRG by incorporating variance reduction techniques [52]. [53] proposed using a Gaussian process to model objective function queries, thereby reducing query complexity and allowing more frequent queries to lower gradient variance. [54] performed random search on a learned low-dimensional manifold, reducing the number of needed objective queries. However, existing ZO accelerators face two main challenges when adapting to ZO fine-tuning for LLMs. First, these approaches were typically designed for smaller-scale tasks involving fewer parameters and less data, and cannot be directly extended to large-scale LLMs. Second, many prior methods focus on improving query efficiency, whereas recent work has shown that a single query can suffice for LLM fine-tuning [14]. How to effectively accelerate ZO optimization on large model fine-tuning remains a problem.
Moreover, ZO has several properties that make it well-suited for on-device or edge training scenarios. 1) Memory efficiency: Edge devices such as mobile phones and FPGAs typically offer limited memory resources. ZO significantly reduces memory usage by avoiding activation and gradient storage, making it more deployable in such constrained setting. 2) Forward-only optimization: As ZO only relies on forward passes, it is compatible with existing inference accelerators (e.g., NNAPI on Android, edge TPUs, etc.), which typically lack support for backpropagation. This makes ZO a strong candidate for adapting inference-only hardware for training.
@@ B.1 Datasets and Evaluation
Table B.1: The hyperparameter for experiments. For DiZO and DiZO LoRA, we only show the setting of extra hyperparameters, and have the same setting in other common hyperparameters with MeZO and MeZO LoRA respectively.
@@
For the RoBERTa-large model, we use the following classification datasets: SST-2 [55], SST-5 [55], SNLI [56], TREC [57], MNLI [58], and RTE [59, 60, 61, 62]. Following previous studies, we cap the test set size at 1000 samples. Two training settings are considered: $k=16$ and $k=512$, where we randomly select 16 or 512 samples per class for both training and validation.
For the OPT and Llama series models, we use the SuperGLUE benchmark [27], which includes RTE [59, 60, 61, 62], CB [63], BoolQ [64], WIC [65], WSC [66], and MultiRC [67]. We also include SST-2 [55] and two question answering datasets, SQuAD [68] and DROP [69]. For each of these datasets, we randomly sample 1000 instances for training, 500 for validation, and 1000 for testing.
B.2 Hyperparameter Setting
We use the hyperparameters in Table B.1 for experiments on RoBERTa-large, OPT-series, and Llama-series models. Specifically, the choice of clip range did not significantly impact the performance. The selection of the projection update cycle and scalar for projection affects the performance somewhat. Generally, for datasets that need larger iterations for convergence, or for these harder datasets, DiZO prefers a larger update cycle, while for those less complicated datasets, DiZO benefits from a smaller update cycle.
Appendix C Ablation study on DiZO
@@ C.1 Ablation for Projection Layers Selection
Instead of applying projections to all layers, which would require storing the entire pre-trained model, we focus only on projecting the weights of the Query and Value in the attention modules. As shown in Table C.1, this strategy achieves the best trade-off between the overall performance and extra storage requirements, does not reduce the performance and only 16.7% of the parameters of the pre-trained model are needed to store. A Similar strategy has also been adopted in LoRA [17].
Table C.1: Ablation study for selecting which layers to project. The highlighted line with a blue rectangle is the setting used in DiZO. Extra memory indicates the extra memory needed due to pre-trained model storing. Attn_Q: attention Query layer; Attn_V: attention Value layer; Attn_K: attention Key layer; Attn_O: attention output projection; Dense: dense fully connected layer.
@@ C.2 Ablation for Strategies in ZO Projection Learning
As discussed in Section 3.3, we introduce two strategies, re-initialization (Re-init) and projection clipping (Clipping), to enhance projection learning and improve the stability of fine-tuning. The ablation results for these strategies, along with the corresponding loss curves, are shown in Figure C.1.
Overall (left in Figure C.1), omitting either Re-init or Clipping significantly diminishes the benefits of DiZO, with MeZO outperforming DiZO in these cases. Specifically, without Re-init, accuracy drops sharply, falling below MeZO. Similarly, without Clipping, while DiZO slightly outperforms MeZO on simpler datasets like SST-2, it suffers from severe model collapse on more challenging datasets, leading to a significant decline in accuracy.
From the loss curve trajectory (right in Figure C.1), without Re-init, DiZO loses its advantage in training acceleration, as the loss curve becomes noticeably slower to decrease. Without Clipping, the loss curve exhibits significant oscillations during certain training steps. This instability arises when projections are optimized to unsuitable values, such as extremely large or small magnitudes. These inappropriate projections cause substantial changes in model weights, leading to pronounced oscillations in the loss.
@@
Figure C.1: Ablation study for the two strategies: re-initialization and projection clipping, which is conducted on RoBERTa-large ($k=16$). Left: overall results when ablating the strategies. Right: loss curve when ablating the strategies.
C.3 Ablation for the Selection of the Anchor
We conduct an experiment to illustrate the effect of other anchor points beyond the pre-trained weights, the results are shown in Table C.2. In conclusion, using $\bm{0}$ or $\bm{\theta_{t-1}}$ significantly reducing the benefit of our method. Specifically, using $\bm{0}$ as an anchor yields similar results to using $\theta_{t-1}$ in terms of GPU hours, but causes unstable training, the accuracy decreases at the later training stage, and causes the results to be even worse than MeZO. Therefore, the effectiveness of our method is inseparable from the choice of anchors. Selecting a more robust anchor could not only improve accuracy but also the convergence speed.
Table C.2: Comparison on conducting projection on learning rate (LR) or use weight at (t1)(t-1)-th iteration $\bm{\theta}_{t-1}$ instead of the weight of the pre-trained model $\bm{\theta_{0}}$ as the base of projection. Results are obtained by fine-tuning OPT-2.7B.
Anchor SST-2 RTE SQuAD
@@ C.3 Ablation for the Selection of the Anchor
C.4 Ablation on Hyperparameter Setting
Figure C.2: Hyperparameter sensitive testing, including clip range, smooth scalar, and update frequency. Evaluated by training Roberta-large on SST2 and SST5 datasets.
We conduct experiments to investigate the effect of different hyperparameter settings, including different clip ranges, smoothing scalars, and projection update frequencies. Results are obtained by fine-tuning RoBERTa-large on SST-2 and SST-5 with 3 different settings for the 3 hyperparameters, as shown in Figure C.2. In conclusion, for easier classification tasks like SST-2, the weight of the model changes fast, and so as the convergence speed, therefore, we can apply more aggressive projection strategies, apply a larger clip range, a larger smoothing scalar, and update the projection more frequently. In contrast, for the rather more difficult tasks, a more conservative projection strategy is preferred.
C.5 Quantization of Anchor (Pre-trained) Model
Our method introduce a anchor model (i.e., the pre-trained model) for projection, roughly 16% of the pretrained parameters must still reside in memory. To further reduce extra memory cost, we explored anchor compression via quantization of the pretrained Query and Value matrices to 8-bit and 4-bit precision [48]. As shown in Table C.3, DiZO can effectively incorporate with the quantization technique, still preserving advantages in both accuracy and GPU hours. Exploring more advanced quantization or compression methods to further improve anchor efficiency while maintaining performance is an important avenue for future work.
Table C.3: DiZO with quantized anchor.
Method SST-2 RTE
Acc GPU hours Acc GPU hours
MeZO 90.0 100% 63.5 100%
DiZO (8-bits) 92.2 63% 67.2 71%
DiZO (4-bits) 91.7 67% 65.2 68%
DiZO 92.5 56% 68.4 62%
Appendix D Do Other Alternative Strategies Work?
As discussed in Section 6.5, we compare with two representative straightforward alternative strategies, Adam and Backtracking linesearch, to highlight the effectiveness of our method.
[…]
Table D.1: Results on fine-tuning OPT-2.7B on SST2 when using Adam as the optimizer. Adam is either memory-intensive or introduces a lot of extra computational overhead.
Method Optimizer Acc. Training FLOPs Memory Iter.
MeZO SGD 85.2 100% 6.8 GB 2K
MeZO Adam (Recompute) 86.1 431% 6.8 GB 2K
MeZO Adam (Store) 86.1 100.2% 17.6GB 2K
DiZO SGD 86.3 61% 7.5GB 1K
DiZO SGD 89.8 122% 7.5GB 2K
Line Search appears to be a simpler alternative for tuning the projection scalar. However, it requires freezing all other layers when optimizing the scalar for a particular layer, resulting in significant computational overhead. To empirically validate this limitation, we replace the ZO-based projection search in DiZO with Armijo-style backtracking line search. For fairness, the line search is initialized at $(1+\tau)\times\text{original learning rate}$, matching the scale used in DiZO. The results in Table D.2 show that, under the same forward pass budget, backtracking performs notably worse than the original MeZO. Furthermore, even when increasing the number of forward passes to 7800, it only matches MeZO’s performance and still falls short of DiZO. These results confirm that traditional line search is not only less effective but also less efficient in the context of high-dimensional, layer-wise projection tuning.
Table D.2: Results on fine-tuning OPT-2.7B on SST2. For fair comparison, we use Armijo-style backtracking line search to replace ZO in our method.
Searching Strategy Forward Pass Accuracy
NA (MeZO) 4000 85.2
Backtracking 4000 81.6
Backtracking 7800 85.1
ZO (DiZO) 4000 89.3
Appendix E More Experimental Results
In this section, we provide a comprehensive presentation of our results across various datasets and models to complement the main paper. Specifically, the results include:
  • Detailed accuracy number and trajectory of training loss on RoBERTa-large (Table E.1 and Figure E.1).
  • More memory and speed results of fine-tuning OPT-2.7B on SST-2 and SQuAD datasets (Table E.2 and Table E.3).
  • Results on larger model, OPT-13B (Table E.4).
  • Results on Llama3-3B and Llama-8B (Table E.5 and Table E.6).
  • Results on more challenging benchmark MMLU and MT-Bench (Table E.7 and Table E.8).
E.1 RoBERTa-large Experiments
Table E.1 reports the corresponding detailed numbers from Figure 2, and Figure E.1 shows the trajectory of training loss.
Table E.1: Experiment results on RoBERTa-large (350M) on six classification datasets. Results of the baseline methods MeZO and MeZO LoRA are taken from [14]. All reported numbers are averaged accuracy with standard deviation shown. Better results between MeZO and DiZO are highlighted in bold.
Dataset Task Type SST-2 SST-5 SNLI MNLI RTE TREC
——-sentiment——- ———-language inference———- –topic–
Zero-shot 79.0 35.5 50.2 48.8 51.4 32.0
Gradient-free methods: $k=16$
MeZO 90.5 (1.2) 45.5 (2.0) 68.5 (3.9) 56.5 (2.5) 59.4 (5.3) 76.9 (2.7)
HiZOO 91.1 (1.6) 46.1 (1.3) 69.3 (3.1) 58.8 (3.1) 57.4 (6.2) 77.2 (2.2)
HiZOO LoRA 91.4 (0.9) 44.8 (1.5) 70.7 (5.2) 64.3 (2.8) 66.9 (3.2) 72.7 (7.3)
DiZO 92.2 (0.9) 47.1 (1.3) 71.0 (3.1) 60.1 (3.5) 67.9 (4.7) 77.4 (2.4)
Gradient-based methods: $k=16$
FT 91.9 (1.8) 47.5 (1.9) 77.5 (2.6) 70.0 (2.3) 66.4 (7.2) 85.0 (2.5)
Gradient-free methods: $k=512$
MeZO 93.3 (0.7) 53.2 (1.4) 83.0 (1.0) 78.3 (0.5) 78.6 (2.0) 94.3 (1.3)
HiZOO 93.5 (0.4) 53.5 (1.2) 83.3 (1.4) 77.2 (1.5) 79.1 (1.2) 94.9 (1.7)
HiZOO LoRA 94.3 (0.5) 54.1 (0.6) 82.7 (1.8) 79.0 (0.8) 80.9 (1.6) 93.1 (0.5)
DiZO 94.6 (0.1) 53.6 (1.7) 84.5 (0.6) 79.8 (0.9) 80.3 (1.8) 93.8 (1.5)
Gradient-based methods: $k=512$
FT 93.9 (0.7) 55.9 (0.9) 88.7 (0.8) 84.4 (0.8) 82.7 (1.4) 97.3 (0.2)
Figure E.1: Trajectory of training loss curves when using MeZO and DiZO to fine-tune Roberta-large on SST-2, MNLI, and RTE.
E.2 More Memory and Speed Analysis
We present the memory and speed results for OPT-2.7B on the SST-2 and SQuAD datasets in Table E.2 and Table E.3, respectively. DiZO significantly reduces the number of required iterations while maintaining throughput comparable to MeZO, leading to substantially fewer training GPU hours. In contrast, HiZOO achieves only modest iteration savings and further reduces the throughput of MeZO by approximately 1.5× due to its reliance on second-order information estimation. As a result, HiZOO offers only a slight improvement over MeZO in terms of training GPU hours. In some cases, such as HiZOO combined with LoRA on SQuAD, it even consumes more training GPU hours than MeZO with LoRA.
@@ E.2 More Memory and Speed Analysis
Table E.2: Memory utilization and speed test on OPT-2.7B on SST-2 dataset (35 tokens per example on average). : partial gradient-free; : gradient-free; : gradient-based. DiZO: searching projection with Adam.
@@
Table E.3: Memory utilization and speed test on OPT-2.7B on SQuAD dataset (300 tokens per example on average). : partial gradient-free; : gradient-free; : gradient-based. DiZO: searching projection with Adam.
@@
E.3 Larger OPT Models Fine-tuning
To further illustrate the generalizability of our method, we conduct experiments on OPT-13B. The results are shown in Table E.4, DiZO consistently outperforms the baselines both in terms of accuracy and speed.
Table E.4: Experiment results on OPT-13B (with 1000 training samples). Better results are highlighted in bold.
SST-2 RTE SQuAD
Dataset Acc GPU hours Acc GPU hours Acc GPU hours
MeZO 91.4 100% 66.1 100% 84.7 100%
HiZOO 92.1 86% 69.3 82% 82.9 91%
DiZO 92.4 69% 72.6 76% 85.2 73%
E.4 Llama Fine-tuning
To demonstrate the generalizability of DiZO, we conducted experiments on the Llama-series models. The results for Llama3-3B and Llama3-8B are presented in Table E.5 and Table E.6, respectively. DiZO consistently outperforms MeZO across both the 3B and 8B Llama models.
However, we observed that ZO LoRA performs poorly with Llama models (including DiZO, MeZO and HiZOO). The loss value remains stagnant, and the resulting accuracy is comparable to or even worse than zero-shot results. We leave it to future work to investigate why ZO LoRA fails with Llama models. We suspect that this limitation may be related to the Group Query Attention (GQA) [70] mechanism employed in Llama3.
Table E.5: Experimental results on Llama3-3B for seven classification datasets and two text generation datasets (with 1000 training samples). Better results between MeZO and DiZO are highlighted in bold.